OpenStudy (kayla2001):

find the first four terms of the sequence represented by the expression. 3n+5 a.)3,8,14,20 b.)8,11,14,17 c.)3,6,9,12 d.)0,8,11,14

OpenStudy (freckles):

assuming the first term is given when n=1... if you replace n with 1; what is given by 3n+5

OpenStudy (kayla2001):

8?

OpenStudy (freckles):

yes 3+5=8 is the first term

OpenStudy (freckles):

to find the 2nd term replace n with 2

OpenStudy (kayla2001):

it would be 11?

OpenStudy (freckles):

right we could keep going but it should be obvious by this point what your answer is

OpenStudy (freckles):

\[3(1)+5,3(2)+5,3(3)+5,3(4)+5,.... \\ = \\ 3+5,6+5,9+5,12+5,... \\ = \\ 8,11,14,17,...\]

OpenStudy (freckles):

just entered in 1,2,3, and then 4 to find the first 4 terms :)